With a tailwind a jet flew 10000 miles In 4 hours the jets return trip against the same tailwind was 5 hours find the jets speed and the wind speed
Answer by mananth(16949)

against wind 5 hours
with wind 4 hours

Distance = same
Jet speed =x
wind speed =y
t=d/r
10000 / ( x - y )= 5.00
5 ( x - y ) = 10,000.00
5 x -5 y = 10000 ....................1
10000 / ( x + y )= 4
4 ( x + y ) = 10000
4 x + 4 y = 10000 ...............2
Multiply (1) by 4
Multiply (2) by 5
we get
20 x -20 y = 40000
20 x + 20 y = 50000
40 x = 90000
/ 40
x = 2250 mph Jet speed

plug value of x in (1)
5 x -5 y = 10000
11250 -5 y = 10000
-5 y = 10000 -11250
-5 y = -1250
y = 250 mph wind speed